A population has a mean of 200 and a standard deviation of 60. Suppose a sample of size 100 is selected and is used to estimate .

What is the probability that the sample mean will be within +/- 6 of the population mean (to 4 decimals)? (Round z value in intermediate calculations to 2 decimal places.)
Answer #1

Solution:- Given that mean = 200, standard deviation = 60,

P(200-6 < X < 200+6) = P(194 < X < 206)

=> P(194 < X < 206) = P((194-200)/60 < (X-μ)/σ < (206-200)/60)

= P(-0.10 < Z < 0.10)

= 0.0796